OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

\[ \|\mathbf u\|\| \mathbf v\|\cos(\theta) = \mathbf u\cdot \mathbf v \]

OpenStudy (anonymous):

In this case \[ \|\mathbf u\| = \sqrt{\mathbf u \cdot \mathbf u} \]

OpenStudy (jhannybean):

sqrt(u * u)? Or.. sqrt[(x-comp)^2 * (y-comp)^2]?

OpenStudy (jhannybean):

\[\|\mathbf u\|\| \mathbf v\|\cos(\theta) = \mathbf u\cdot \mathbf v\]\[\|\mathbf u\|\ = \sqrt{(-5)^2 +(-4)^2} = \]\[\|\mathbf v\|\ = \sqrt{(-4)^2+(-3)^2}= \]\[\mathbf u\ \cdot \mathbf v = \]\[\cos \theta = \frac{(\|\mathbf u\| \cdot \|\mathbf v\|\ )}{\|\mathbf u\|\ \| \mathbf v\|}=\]